OpenStudy (anonymous):

2sin(3x)-square root 2=0

OpenStudy (mathmale):

although I could guess what the objective of this problem is, I'd appreciate your including the instructions.

OpenStudy (anonymous):

solve for x thanks

OpenStudy (anonymous):

sin 3x = (sqr.2)/2 = sin Pi/4 = sin 3Pi/4 a) 3x = Pi/4 --> x = Pi/12 b) 3x = 3Pi/4 --> x = Pi/4

OpenStudy (mathmale):

@thu1935 : Thank you for your contribution. Please guide the questioner through the solution of the problem so that he/she may find his/her own answers. Do not give away the answers.

OpenStudy (mathmale):

@mparent : Could you possibly use the Equation Editor to present your math problem?\[2\sin(3x)-\sqrt{2}=0\] What would you do next?